The equation y=-1777x+27,153 can be used to predict the number y of gun deaths in the United States x years after 2000, that is x=0 corresponds to 2000,x=3 corresponds to 2003 x=4 corresponds to 2004, and so on. Predict the number of gun deaths in 2004 and 2009. In what year will the number of gun deaths be 7606?
The predicted gun deaths in 2004 will be _____.
The predicted gun deaths in 2009 will be ______.
The predicted gun deaths in 7606 in the year _____.

Found 2 solutions by checkley77, stanbon:
Answer by checkley77(12844)

y=-1777x+27,153
2004=-1,777*4+27,153
2004=-7,108+27,153
2004=20,065 ans.
2009=-1,777*9+27,153
2009=-15,993+27,173
2009=11,180 ans
7,606=-1,777x+27,173
7,606=-1,777x+27,173
1,777x=-7,606+27,173
1,777x=19,567
x=19,567/1,777
x=11.01 or the year 2011 the gun death will be 7,606.
Proof:
2011=-1,777*11.01+27,153
2011=-19,564.77+27,153
2011~7,606
You also asked to predict the number of deaths in the year 7606.
7606=-1,777*5,606+27,173
7606=-9,961,862+27,173
7606=-9,934,689 deaths. Not a valid ans.

Answer by stanbon(75887)

The equation y=-1777x+27,153 can be used to predict the number y of gun deaths in the United States x years after 2000, that is x=0 corresponds to 2000,x=3 corresponds to 2003 x=4 corresponds to 2004, and so on. Predict the number of gun deaths in 2004 and 2009. In what year will the number of gun deaths be 7606?
The predicted gun deaths in 2004 will be 1777*4 + 27,153 _____.
The predicted gun deaths in 2009 will be 1777*9 + 27,153 ______.
The predicted gun deaths in 7606 in the year 1777*? + 27,153 _____.
------
Did you really mean 7606 which is 5606 after the year 2000?
===================
Cheers,
Stan H.